OpenStudy (anonymous):

solve for x in following equation pls x+1/x+6 = x-1/x=3

OpenStudy (anonymous):

solve for x in following equation pls x+1/x+6 = x-1/x=3

OpenStudy (anonymous):

Are you sure its x=3 as the denominator on the second half of the equation? or meant to be x+3?

OpenStudy (anonymous):

it is x+3 but the problem has been answered can you help with this one plse: 2-a-4/a+3 = a^2-2/a+3

jhonyy9 (jhonyy9):

- now i have resolved it - check please where you have uploaded it again

jhonyy9 (jhonyy9):

this equation with a